OpenStudy (anonymous):

The width of a rectangle increases at 2 cm/s, while the length decreases at 3cm/s. How fast is the area of the rectangle changing when the width equals 20 cm and the length equals 50cm?

OpenStudy (amistre64):

A = l*w dA = l dw + w dl dw = 2 ; dl = -3

OpenStudy (amistre64):

dA = 50(2) + 20(-3) dA = 100 -60 = 40 cm^2 / sec

OpenStudy (amistre64):

a square is the biggest area, so until it reaches that point, it increases i think